# RankMamba: Benchmarking Mamba’s Document Ranking Performance in the Era of Transformers

### v2

@@ 4 Result and Analysis
Document Ranking Performance. We report the ranking performance in Table 3. We notice that encoder-only language models such as BERT  and RoBERTa  can achieve strong performance. For example, within models $\approx$ 110M and models $\approx$ 330M parameters, bert-base-uncased and roberta-large achieve the highest MRR on MSMARCO Dev dataset.
@@
With the current training recipe, Mamba  models achieve competitive performance compared to transformer-based models of similar scales. For example, among models $\approx$ 110M parameters, state-spaces/mamba-130m achieves the highest NDCG@10 on TREC DL19 and DL20 datasets. state-spaces/mamba-370m also achieves highest NDCG@10 on DL19, and second highest NDCG@10 on DL20 among 5 models we compare with scale of $\approx$ 330M parameters. Mamba’s strong performance is consistent when trained with LoRA. For example, state-spaces/mamba-790m compared to T5-large and EleutherAI/pythia-1b, approaching full parameter fine-tuning of state-spaces/mamba-370m.
Benchmarking Training Throughput for Document Ranking. We benchmark the training throughput of different models on a server with one A6000 GPU and report results in Figure 1. Overall, models trained with flash attention implementation (OPT, Pythia) achieve the highest training throughput and lowest space complexity. We notice that although Mamba  can achieve $O(n)$ training complexity in theory, the training throughput with its current implementation is still lower than efficient transformer implementations such as flash attention. This finding suggests the deficiency of Mamba’s current implementation and leaves us with potential future work for improvement.
Figure 1: We show the training throughput of models $\approx$330M parameters and models $>$ 700M parameters. Models $>$ 700M parameters are trained with LoRA, with rank=32. We notice Mamba  models have lower throughput and higher GPU memory consumption compared to efficient transformer implementation such as flash attention.
@@ 5 Related Works
Transformer-based Language Models utilize the attention mechanism [Vas+17] to produce contextualized representations of tokens and demonstrate superior performance and training efficiency compared to traditional recurrent neural network models [HS97, Pet+18]. Such LMs can be used as feature extractors for downstream tasks or be used for natural language generation (NLG) with a specific language modeling head. Prior works [XJ24, Raf+20, Wan+22, Le ̵+22] have shown that between models of similar sizes, different pre-training objectives can lead to strengths and weaknesses in terms of natural language understanding (NLU) and natural language generation capabilities.
Language Models in IR. Neural network-based language modeling techniques have gained widespread adoption within the information retrieval (IR) community, including web search snippet generation [BC14, XC23, Xu23], modeling users and products in recommender systems [XZA21, ZXA21], generating search results explanation [YRA22, Xu+23] and conversational search [Wan+23, WXA23, Ali+19].
Document Ranking Task aims to capture the relevance between query and document. Traditional bag-of-words methods such as BM25 [RZ+09] aim to capture the lexical overlapping. Neural models such as word2vec [Mik+13] are learned from static word co-occurrence patterns from a pre-training corpus and can capture the semantic meanings of the words. Transformer-based language models can capture contextualized representations and have been widely applied to modeling query-document relevance. [DC19, ] propose to model long documents with BERT [Dev+19]. To model long documents with BERT’s limited 512 context window, they propose several model variants such as FirstP, MaxP and SumP. Followup works [Li+23, Hof+21, Boy+22, Ma+23] achieve performance improvements by leveraging techniques such as mining hard negatives, chunking long documents and/or sliding windows, and replacing BERT with better, long-context pre-trained language models such as Longformer [BPC20] and Llama [Tou+23]. For a more comprehensive literature review, refer to [LNY22, ].
@@ 6 Conclusion and Limitations
This study aims to benchmark Mamba  models’ performance in the document ranking task. We find that Mamba  models can achieve competitive performance compared to transformer-based language models, and the effectiveness is consistent with trained with LoRA. However, Mamba  models have the deficiency of lower training throughput compared to efficient attention implementations, limiting their potential for efficient training and deployment.
Due to the limited bandwidth, we only study models $<$ 1B parameters, our current finding may change when scaling to larger models and adopting different training recipes. The efficacy of Mamba  models for other classical IR tasks such as document retrieval remains to be studied.

### v3

@@ 4 Result and Analysis
Document Ranking Performance. We report the ranking performance in Table 3. We notice that encoder-only language models such as BERT and RoBERTa can achieve strong performance. For example, within models 110M and models 330M parameters, bert-base-uncased and roberta-large achieve the highest MRR on MSMARCO Dev dataset.
@@
With the current training recipe, Mamba models achieve competitive performance compared to transformer-based models of similar scales. For example, among models 110M parameters, state-spaces/mamba-130m achieves the highest NDCG@10 on TREC DL19 and DL20 datasets. state-spaces/mamba-370m also achieves highest NDCG@10 on DL19, and second highest NDCG@10 on DL20 among 5 models we compare with scale of 330M parameters. Mamba’s strong performance is consistent when trained with LoRA. For example, state-spaces/mamba-790m compared to T5-large and EleutherAI/pythia-1b, approaching full parameter fine-tuning of state-spaces/mamba-370m.
Benchmarking Training Throughput for Document Ranking.
We benchmark the training throughput of different models on a server with one A6000 GPU and report results in Figure 1. Overall, models trained with flash attention implementation (OPT, Pythia) achieve the highest training throughput and lowest space complexity. We notice that although Mamba can achieve $O(n)$ training complexity in theory, the training throughput with its current implementation is still lower than efficient transformer implementations such as flash attention. This finding suggests the deficiency of Mamba’s current implementation and leaves us with potential future work for improvement.
Figure 1: We show the training throughput of models 330M parameters and models >> 700M parameters. Models >> 700M parameters are trained with LoRA, with rank=32. We notice Mamba models have lower throughput and higher GPU memory consumption compared to efficient transformer implementations such as flash attention.
@@ 5 Related Works
Transformer-based Language Models
These language models utilize the attention mechanism (Vaswani et al., 2017) to produce contextualized representations of tokens and demonstrate superior performance and training efficiency compared to traditional recurrent neural network models (Hochreiter and Schmidhuber, 1997; Peters et al., 2018). Such LMs can be used as feature extractors for downstream tasks or be used for natural language generation (NLG) with a specific language modeling head. Prior works (Xu and Jiang, 2024; Raffel et al., 2020; Wang et al., 2022; Le Scao et al., 2022) have shown that between models of similar sizes, different pre-training objectives can lead to strengths and weaknesses in terms of natural language understanding (NLU) and natural language generation capabilities.
Language Models in IR
Neural network-based language modeling techniques have gained widespread adoption within the information retrieval (IR) community, including web search snippet generation (Bast and Celikik, 2014; Xu and Cohen, 2023; Xu, 2023), modeling users and products in recommender systems (Xu et al., 2021; Zeng et al., 2021), generating search results explanation (Yu et al., 2022; Xu et al., 2023) and conversational search (Wang et al., 2023b, a; Aliannejadi et al., 2019).
Document Ranking
The document ranking task aims to capture the relevance between query and document. Traditional bag-of-words methods such as BM25 (Robertson et al., 1999) aim to capture the lexical overlapping. Neural models such as word2vec (Mikolov et al., 2013) are learned from static word co-occurrence patterns from a pre-training corpus and can capture the semantic meanings of the words. Transformer-based language models can capture contextualized representations and have been widely applied to modeling query-document relevance. Dai and Callan (2019) propose to model long documents with BERT (Devlin et al., 2019). To model long documents with BERT’s limited 512 context window, they propose several model variants such as FirstP, MaxP and SumP. Followup works (Li et al., 2023; Hofstätter et al., 2021; Boytsov et al., 2022; Ma et al., 2023) achieve performance improvements by leveraging techniques such as mining hard negatives, chunking long documents and/or sliding windows, and replacing BERT with better, long-context pre-trained language models such as Longformer (Beltagy et al., 2020) and Llama (Touvron et al., 2023a, b). For a more comprehensive literature review, refer to Lin et al. (2022) and Xu et al. (2025c)
@@ 6 Conclusion and Limitations
This study aims to benchmark Mamba models’ performance in the document ranking task. We find that Mamba models can achieve competitive performance compared to transformer-based language models, and the effectiveness is consistent with trained with LoRA. However, Mamba models have the deficiency of lower training throughput compared to efficient attention implementations, limiting their potential for efficient training and deployment. We only explored the contrastive learning training strategy. As recent works demonstrate, knowledge distillation is especially effective for training lightweight rerankers when distilling from a more capable reranker teacher (Baldelli et al., 2024; Schlatt et al., 2025; Xu et al., 2025b). Due to the limited bandwidth, we only study reranker models << 1B parameters based on the BM25 retriever. Our future work should include extending to rerankers of large scales, different document retrievers.